OpenStudy (anonymous):

Plot the y-intercept and the vertex of the following function: g(x) = - x2 + 6x - 8 What is the smallest zero of the function?

hero (hero):

g(x) = - x2 + 6x - 8 0 = -x^2 + 6x - 8 0 = -(x^2 - 6x + 8) 0 = x^2 - 6x + 8 0 = x^2 -4x - 2x + 8 0 = x(x-4)-2(x-4) 0 = (x-4)(x-2) x = 2,4 Therefore, 2 is the smallest zero.

OpenStudy (anonymous):

thank you! (:

hero (hero):

yw :D